OpenStudy (anonymous):

determine whether the graphs of the equations are perpendicular 4x-3y=2 and 4y-3x=5

OpenStudy (anonymous):

determine whether the graphs of the equations are perpendicular 4x-3y=2 and 4y-3x=5

OpenStudy (anonymous):

get both in y=mx +b form , then find the gradients

OpenStudy (anonymous):

first one y= (2-4x) / -3 , so the gradient is -4/-3 = 4/3 second one, y= (5+3x) / 4 , so the gradient is 3/4

OpenStudy (anonymous):

for these to be perpendicular we need the product of the gradient to be -1. In this case the product of the graident is +1 , so they are not perpendiculat

OpenStudy (anonymous):

perpendicular*

OpenStudy (anonymous):

thanks elecengineer i appreciate you